OpenStudy (abbyrose0105):

Divide x to the 1 half power divided by x to the 2 sevenths power.

OpenStudy (abbyrose0105):



OpenStudy (campbell_st):

well dividing, the index law is subtract the powers... so its \[x^{\frac{1}{2} - \frac{2}{7}} = x^?\] find the value of ?

OpenStudy (abbyrose0105):

So it would be x^3/14

OpenStudy (jdoe0001):

\(\Large { \cfrac{1}{\frac{1}{a^{\color{red} n}}}\implies a^{{\color{red} n}} \\ \quad \\ \quad \\ \cfrac{x^{\frac{1}{2}}}{x^{\frac{2}{7}}}\implies \cfrac{x^{\frac{1}{2}}}{1}\cdot \cfrac{1}{x^{\frac{2}{7}}}\implies x^{\frac{1}{2}}\cdot x^{-\frac{2}{7}}\implies ? }\)

OpenStudy (abbyrose0105):

I got it right on my exam. Thanks!

OpenStudy (jdoe0001):

hmmm I even got the wrong.... set... anyhow \(\Large \bf { \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}}\implies a^{-\frac{{\color{blue} n}}{{\color{red} m}}} \\ \quad \\ \quad \\ \cfrac{x^{\frac{1}{2}}}{x^{\frac{2}{7}}}\implies \cfrac{x^{\frac{1}{2}}}{1}\cdot \cfrac{1}{x^{\frac{2}{7}}}\implies x^{\frac{1}{2}}\cdot x^{-\frac{2}{7}}\implies ? }\)

OpenStudy (abbyrose0105):

My answer was x^3/14

OpenStudy (jdoe0001):

yeap